# S-DAG: A Subject-Based Directed Acyclic Graph for Multi-Agent Heterogeneous Reasoning

### v1

@@ Abstract
Large Language Models (LLMs) have achieved impressive performance in complex reasoning problems. Their effectiveness highly depends on the specific nature of the task, especially the required domain knowledge. Existing approaches, such as mixture-of-experts, typically operate at the task level; they are too coarse to effectively solve the heterogeneous problems involving multiple subjects. This work proposes a novel framework that performs fine-grained analysis at subject level equipped with a designated multi-agent collaboration strategy for addressing heterogeneous problem reasoning. Specifically, given an input query, we first employ a Graph Neural Network to identify the relevant subjects and infer their interdependencies to generate an Subject-based Directed Acyclic Graph (S-DAG), where nodes represent subjects and edges encode information flow. Then we profile the LLM models by assigning each model a subject-specific expertise score, and select the top-performing one for matching corresponding subject of the S-DAG. Such subject-model matching enables graph-structured multi-agent collaboration where information flows from the starting model to the ending model over S-DAG. We curate and release multi-subject subsets of standard benchmarks (MMLU-Pro, GPQA, MedMCQA) to better reflect complex, real-world reasoning tasks. Extensive experiments show that our approach significantly outperforms existing task-level model selection and multi-agent collaboration baselines in accuracy and efficiency. These results highlight the effectiveness of subject-aware reasoning and structured collaboration in addressing complex and multi-subject problems.
@@ 1 Introduction
In recent years, intelligent agents based on large language models (LLMs) have developed rapidly and achieved significant advancements across various fields, ranging from question answering (yue2025survey; zhuang2023toolqa) to text generation (huang2023grounded; wu2024autogen) and complex reasoning tasks (ke2025survey; zhang2024llm). While the development of a general-purpose LLM is promising (mumuni2025large; kojima2022large), it becomes evident that a single LLM often struggles to handle complex reasoning tasks, especially when these problems span multiple disciplines (feng2025one). This limitation raises higher demands for model training and fine-tuning (hoffmann2022training). In this context, multi-agent systems based on LLMs have emerged (guo2024large; du2023improving; talebirad2023multi; han2024llm; gu2025explain; liang2024encouraging), aiming to leverage the collective intelligence and specialized expertise of multiple agents to tackle complex and multidisciplinary problems [^1] For simplicity, agents, LLMs, and models are used interchangeably..
@@ 2 Related Work
Multi-Agent Systems. Recent advances in Multi-Agent Systems (MAS) have introduced diverse collaboration mechanisms to tackle complex tasks (guo2024large; talebirad2023multi; han2024llm), broadly categorized into fixed and dynamic paradigms. (1) Fixed Multi-Agent Systems rely on manually designed architectures, such as LLM debates (du2023improving; liang2024encouraging), chain-of-agents (gu2025explain; zhang2024chain; tao2025chain), and graph-based systems (yin2023exchange; li2024improving). These approaches enable collaboration among agents with predefined roles and structures, making them effective for well-structured problems, but often lacking adaptability to dynamic tasks. (2) Dynamic Multi-Agent Systems adapt their structure in response to real-time task demands. Notable examples include GPTSwarm (zhuge2024gptswarm) and Heterogeneous Swarms (feng2025heterogeneous), which optimize collaboration through dynamic graph structures. AgentPrune (zhang2024cut) improves communication efficiency by pruning redundant links, while DyLAN (liu2024dynamic) dynamically selects agents and communication paths based on task context. MasRouter (yue2025masrouter) introduces a cascaded controller for mode selection, role assignment, and LLM routing, enabling efficient and adaptive MAS construction.
@@ 4.1 Experiment Setup
Our selection of baselines is guided by the goal of evaluating the challenge of heterogeneous reasoning, where a single complex question spans multiple subject domains. We consider two primary categories. First, we evaluate single-model methods to test whether a single, high-performance, general-purpose LLM can effectively handle multi-subject reasoning. Specifically, we include models such as closed-source GPT4o-mini (hurst2024gpt), open-source Qwen2.5-72B (yang2024qwen2) and Llama3.3-70B (grattafiori2024llama). Then, the Mixture-of-Experts (MoE) approaches that dynamically select one expert model; GraphRouter that utilize a GNN to select expert model based on contextual information given a problem. Second, we explore multi-model methods, which leverage specialized expertise through collaborative or modular strategies. These include Multi-Agent Debate (MAD), where multiple agents reason through dialogue (liang2024encouraging); and Symbolic-MoE (chen2025symbolic), which conducts skill-level expert selection. These baselines provide a comprehensive framework to assess both generalist and specialist strategies for tackling complex, interdisciplinary reasoning tasks. Prompts for different baselines are shown in Table 8 in appendix.
@@ 4.3 Ablation Study
To quantify the contribution of individual components in our S-DAG framework, we conduct a structured ablation study, as presented in Table 4. This analysis isolates three key design factors: (1) the presence or absence of GNN-based coordination, (2) the use of subject-aware model selection via LLM profiling versus random assignment, and (3) the impact of graph topology, comparing our sparse S-DAG structure to a fully connected alternative. We evaluate each configuration across two core dimensions: task performance and computational efficiency. Inference time denotes the average wall-clock latency required to process a single multi-subject question, encompassing decomposition, agent invocation, message passing, and response synthesis. LLM call count indicates the average number of distinct language model calls per instance, serving as a proxy for computational and monetary cost. More information is detailed in appendix.

### v2

@@ Abstract
Large Language Models (LLMs) have achieved impressive performance in complex reasoning problems. Their effectiveness highly depends on the specific nature of the task, especially the required domain knowledge. Existing approaches, such as mixture-of-experts, typically operate at the task level; they are too coarse to effectively solve the heterogeneous problems involving multiple subjects. This work proposes a novel framework that performs fine-grained analysis at the subject level equipped with a designated multi-agent collaboration strategy for addressing heterogeneous problem reasoning. Specifically, given an input query, we first employ a Graph Neural Network to identify the relevant subjects and infer their interdependencies to generate an Subject-based Directed Acyclic Graph (S-DAG), where nodes represent subjects and edges encode information flow. Then, we profile the LLM models by assigning each model a subject-specific expertise score, and select the top-performing one for matching the corresponding subject of the S-DAG. Such subject-model matching enables graph-structured multi-agent collaboration where information flows from the starting model to the ending model over S-DAG. We curate and release multi-subject subsets of standard benchmarks (MMLU-Pro, GPQA, MedMCQA) to better reflect complex, real-world reasoning tasks. Extensive experiments show that our approach significantly outperforms existing task-level model selection and multi-agent collaboration baselines in accuracy and efficiency. These results highlight the effectiveness of subject-aware reasoning and structured collaboration in addressing complex and multi-subject problems.
Codehttps://github.com/WanyuGroup/AAAI2026˙S-DAG
Extended versionhttps://arxiv.org/abs/2511.06727
@@ 1 Introduction
In recent years, intelligent agents based on large language models (LLMs) have developed rapidly and achieved significant advancements across various fields, ranging from question answering (yue2025survey; zhuang2023toolqa) to text generation (huang2023grounded; wu2024autogen) and complex reasoning tasks (ke2025survey; zhang2024llm). While the development of a general-purpose LLM is promising (mumuni2025large; kojima2022large), it becomes evident that a single LLM often struggles to handle complex reasoning tasks, especially when these problems span multiple disciplines (feng2025one). This limitation raises higher demands for model training and fine-tuning (hoffmann2022training). In this context, multi-agent systems based on LLMs have emerged, aiming to leverage the collective intelligence and specialized expertise of multiple agents to tackle complex and multidisciplinary problems (guo2024large; du2023improving; talebirad2023multi; han2024llm; gu2025explain; liang2024encouraging; yao2025socialized) [^1] For simplicity, agents, LLMs, and models are used interchangeably..
@@ 2 Related Work
Multi-Agent Systems. Recent advances in Multi-Agent Systems (MAS) have introduced diverse collaboration mechanisms to tackle complex tasks (guo2024large; talebirad2023multi; han2024llm), broadly categorized into fixed and dynamic paradigms. (1) Fixed Multi-Agent Systems rely on manually designed architectures, such as LLM debates (du2023improving; liang2024encouraging), chain-of-agents (gu2025explain; zhang2024chain; tao2025chain), and graph-based systems (yin2023exchange; li2024improving; lin2025graph). These approaches enable collaboration among agents with predefined roles and structures, making them effective for well-structured problems, but often lacking adaptability to dynamic tasks. (2) Dynamic Multi-Agent Systems adapt their structure in response to real-time task demands. Notable examples include GPTSwarm (zhuge2024gptswarm), GHG (li2025graphs) and Heterogeneous Swarms (feng2025heterogeneous), which optimize collaboration through dynamic graph structures. AgentPrune (zhang2024cut) improves communication efficiency by pruning redundant links, while DyLAN (liu2024dynamic) dynamically selects agents and communication paths based on task context. MasRouter (yue2025masrouter) introduces a cascaded controller for mode selection, role assignment, and LLM routing, enabling efficient and adaptive MAS construction.
@@ 4.1 Experiment Setup
Our selection of baselines is guided by the goal of evaluating the challenge of heterogeneous reasoning, where a single complex question spans multiple subject domains. We consider two primary categories. First, we evaluate single-model methods to test whether a single, high-performance, general-purpose LLM can effectively handle multi-subject reasoning. Specifically, we include models such as closed-source GPT4o-mini (hurst2024gpt), open-source Qwen2.5-72B (yang2024qwen2) and Llama3.3-70B (grattafiori2024llama). Then, the Mixture-of-Experts (MoE) approaches that dynamically select one expert model; GraphRouter that utilize a GNN to select expert model based on contextual information given a problem. Second, we explore multi-model methods, which leverage specialized expertise through collaborative or modular strategies. These include Multi-Agent Debate (MAD), where multiple agents reason through dialogue (liang2024encouraging); and Symbolic-MoE (chen2025symbolic), which conducts skill-level expert selection. These baselines provide a comprehensive framework to assess both generalist and specialist strategies for tackling complex, interdisciplinary reasoning tasks.
@@ 4.3 Ablation Study
To quantify the contribution of individual components in our S-DAG framework, we conduct a structured ablation study, as presented in Table 4. This analysis isolates three key design factors: (1) the presence or absence of GNN-based coordination, (2) the use of subject-aware model selection via LLM profiling versus random assignment, and (3) the impact of graph topology, comparing our sparse S-DAG structure to a fully connected alternative. We evaluate each configuration across two core dimensions: task performance and computational efficiency. Inference time denotes the average wall-clock latency required to process a single multi-subject question, encompassing decomposition, agent invocation, message passing, and response synthesis. LLM call count indicates the average number of distinct language model calls per instance, serving as a proxy for computational and monetary cost.
@@ 5 Conclusion
6 Acknowledgments
This research was supported by Project P0049179 under the Innovation and Technology Fund – Guangdong–Hong Kong Technology Cooperation Funding Scheme (ITF-TCFS), funded by the Innovation and Technology Commission (Funding Body Ref. No. GHP/386/23SZ).